OpenStudy (shesolitt):

Write the equation of the line that passes through (3, –2) and has a slope of 4 in point-slope form.

OpenStudy (shesolitt):

y + 2 = 4(x – 3) y – 3 = 4(x + 2) x – 3 = 4(y + 2) x + 2 = 4(y – 3)

OpenStudy (shesolitt):

help ! ill medal and fan !

OpenStudy (shesolitt):

@Directrix

OpenStudy (shesolitt):

@leahhhmorgannn

OpenStudy (triciaal):